OpenStudy (anonymous):

find y for 2x+2y=2 and x+2y=1 ?????

OpenStudy (anonymous):

2x+2y=2____A x+2y=1____B A-B---->x=1 2y=0 y=0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can u show the workinng out plz

OpenStudy (anonymous):

can i know how to go to next line while typing the answer

OpenStudy (anonymous):

enter

OpenStudy (anonymous):

hmmmmmmm not gettin:(

OpenStudy (anonymous):

2x+2y =2........... 1 equ x+2y =1..............2 equ 1implies......y=1-x wholediv2 substituing y value 2x+2(1-x/2)=2 2 get canceled 2x+1-x =2 x+1-2=0 x-1=0 substitute x=1in y y=1-1/2 y=0/2 y =0